OpenStudy (anonymous):

Rewrite each angle in degree measure. (Do not use a calculator.) a. 15pi/6 b. 28pi/15

OpenStudy (anonymous):

remember, we treat pi as 180. Rewrite pi as 180 and find the number

OpenStudy (anonymous):

so a. 15(180)/6 b. 28(180)/15 ??

OpenStudy (anonymous):

Yes. The end number will be in degrees.

OpenStudy (anonymous):

in that case I got a. 450 and b. 336

OpenStudy (anonymous):

so it would be 450 degrees and 336 degrees?

OpenStudy (anonymous):

Anything above 360 is equal to an angle below 360 as a reference angle

OpenStudy (anonymous):

what? now I'm confused :o

OpenStudy (anonymous):

450 = 360 + 90 So: 450 degrees is the same as 90 degrees and one complete rotation. E.g.m sin(450 degrees) = sin(90 degrees)

OpenStudy (anonymous):

so just to be clear, A would be sin(450) degrees = sin(90) degrees?

OpenStudy (anonymous):